OpenStudy (anonymous):

I am asking for help here...... It says write a expression in radical notation.... 625-1/4

OpenStudy (anonymous):

-1/4 is the exponent

OpenStudy (anonymous):

\[\frac{1}{\sqrt[4]{625}}\]

OpenStudy (anonymous):

In general: \[y^{-x} = \frac{1}{y^{x}}\] And: \[y^{\frac{1}{x}} = \sqrt[x]{y}\]

OpenStudy (zehanz):

Hey Sabrena, did you notice that 625 = 5^4?...this means that after doing the chore of writing the number in radical notation, as has been done above, you can simplify it to a really simple number!